OpenStudy (anonymous):

9x^2+9x^3-7-13x/3x+4

OpenStudy (anonymous):

3x+4)9x^3+9x^2-13x-7(3x^2-x-1/3 9x^3+12x^2 ----------------------- -3x^2-13x-7 -3x^2-12x --------------------- -x-7 -x-4/3 ------------------------- -17/3

OpenStudy (anonymous):

here 3x^2-x-1/3 is the quotient and -17/3 is remainder

OpenStudy (anonymous):

\[(3x ^{2}-3x-1/3) -17/(9x+12) \]

OpenStudy (anonymous):

\[\frac{9x^2+9x^3-7-13x}{3x+4}=\frac{-9x^3+10x-7}{4}\]

OpenStudy (anonymous):

how Craigwillium?

OpenStudy (anonymous):

can u please show the steps...

OpenStudy (anonymous):

It's a matter of simplification by moving around exponents. I only did the simplification part.

OpenStudy (anonymous):

@craig Oh even i em interested in seeing the proof